This quesion is from aplus.cmcss.net not a book.

max is 24 years older than his son liam. in two years liam will be half as old as max. how old is liam now? and max?
i assumed it could be put like this but it has not been working for me.
m= l + 24
l= 1/2m + 2

Answer by ankor@dixie-net.com(12704)

max is 24 years older than his son liam.
in two years liam will be half as old as max.
how old is liam now? and max?
i assumed it could be put like this but it has not been working for me.
m = l + 24
:
Both will be two years older, so you have
l+2 = .5(m+2)
l + 2 = .5m + 1
l = .5m + 1 - 2
l = .5m - 1
from the 1st equation substitute (l+24) for m
l = .5(l+24) - 1
l = .5l + 12 - 1
l - .5l = 11
.5l = 11
Multiply both sides by 2 to get 1l
l = 22
then
m = 22 + 24
m = 46
:
:
See if that is true in the statement:
"in two years liam will be half as old as max."
22 + 2 = .5(46 + 2)
:
Did this make sense to you now?